OpenStudy (anonymous):

factor completely a^ 2 - a - 20

OpenStudy (anonymous):

1*-20=-20 5-4=1 5*-4=-20 a^2-(5-4)a-20 =a^2-5a+4a-20 now you can proceed further.

OpenStudy (anonymous):

awesome thxs

OpenStudy (anonymous):

yw